Circular No. 7117

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVAE A. Kim, Collège de France, and N. Regnault, Laboratoire de l’Accélér-

ateur Linéaire, Université de Paris-Sud, on behalf of the EROS collaboration; and P. Nugent, Lawrence Berkeley National Laboratory (LBNL), on behalf of the Supernova Cosmology Project, report the discovery of nine supernovae, discovered in the course of the automated supernova search with the 1-m Marly telescope (+ CCD) at La Silla (limiting mag ∼ 21.5). Find-

ing charts can be found at http://www.lal.in2p3.fr/EROS/SNe/Charts/.

SN  1998 UT  α_{2000}  δ_{2000}  V  Offset
1999ad Feb. 9 11^{h}33^{m}18.^{s}46 − 5^{o}45′29′′.8 20.2 0′′.9 W, 2′′.2 S
1999ae Feb. 10 11 51 24.48 − 4 39 09.4 21.2 4′′.6 E, 4′′.6 S
1999af Feb. 12 13 44 50.95 − 6 40 12.6 20.6 3′′.4 E, 4′′.2 N
1999ag Feb. 12 12 15 22.81 − 5 18 12.4 20.7 2′′.1 E, 1′′.7 S
1999ah Feb. 13 12 09 37.20 − 6 18 34.3 20.8 0′′.6 E, 0′′.2 S
1999ai Feb. 15 13 14 10.57 − 5 35 43.7 18.1 14′′.0 E 11′′.3 N
1999aj Feb. 17 11 22 39.34 −11 43 53.9 20.9 1′′.4 W, 4′′.3 S
1999ak Feb. 17 11 06 52.05 −11 39 13.3 18.7 7′′.6 E, 2′′.7 N
1999al Feb. 21 11 10 25.68 − 7 26 37.0 19.4 6′′.4 E, 2′′.7 N

SN 1999ad was confirmed on CCD frames taken on Feb. 12 but was not present on a frame taken on Jan. 13; SN 1999aj was confirmed on CCD frames taken on Feb. 18 but was not present on a frame taken on Jan.

22. SN 1999ae was confirmed with a CCD spectrum taken at the APO 3.5-m telescope by X. Fan (Princeton) and H. Newberg (Fermi National Accelerator Laboratory) on Feb. 17; Nugent, G. Aldering, and S. Deustua (LBNL) note that the spectrum is consistent with a type-II supernova, < 1 month old; z = 0.076. SN 1999af was confirmed with a CCD spectrum taken at the Lick 3.0-m telescope by Aldering on Feb. 25; Nugent and Deustua find that the spectrum is consistent with a type-Ia event, 1 week after peak brightness; z = 0.097. A. Goobar (Stockholm University) and I. Hook (Royal Observatory, Edinburgh) obtained a CCD spectrum of SN 1999ag on Feb. 13 at the Nordic Optical Telescope (NOT); they, along with Nugent and Aldering, find that the host galaxy has z = 0.0988 from O III, Hα, and Hβ lines. The spectrum of SN 1999ag is most consistent with an early type-II supernova at the same redshift; however, the spectrum’s signal-to-noise is too poor to state this with much confidence.

1999 March 3  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green